# Structure of the free interfaces near triple junction singularities in harmonic maps and optimal partition problems

### v1

@@ 1.1. Notations and known results
are open and well-defined.
Before stating our main results in Section 1.2, from Section 1.1.1 to Section 1.1.6 we review some known facts concerning critical points $u\in\mathcal{S}(D;N)$. Most of the properties we recall were proved in a series of papers, namely [GS92, Sun03, CL07, Alp20] for the class $\mathcal{M}(D;N)$ and [CTV05c, TT12] for $\mathcal{S}(D;N)$.
1.1.1. Almgren’s frequency function
First of all, for any non-trivial $u\in\mathcal{S}(D;N)$ one can define the Almgren frequency function for any $x_{0}\in D$ and any $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$ as
$\mathcal{N}(u,r,x_{0}):=\frac{E(u,r,x_{0})}{H(u,r,x_{0})},$
@@ 1.1.1. Almgren’s frequency function
$E(u,r,x_{0}):=\frac{1}{r^{d-2}}E(u,B_{r}(x_{0}))=\frac{1}{r^{d-2}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}|\nabla u_{i}|^{2}\,\mathrm{d}x$
@@
$H(u,r,x_{0}):=\frac{1}{r^{d-1}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}u_{i}^{2}\,\mathrm{d}S$
is the scaled height function. The crucial fact is that the $r\mapsto\mathcal{N}(u,r,x_{0})$ is monotone non-decreasing and so the frequency of $u$ at the point $x_{0}$
$\gamma(u,x_{0}):=\lim_{r\to 0^{+}}\mathcal{N}(u,r,x_{0})$
is well-defined. As a consequence, one can derive unique continuation properties implying that the nodal set
$\mathcal{Z}(u)=\{x\in D\colon u_{i}(x)=0~{}\text{for all }i=1,\dots,N\}$
has empty interior and coincides with the free interface, i.e. $\mathcal{Z}(u)=\mathcal{F}(u)$. Moreover, there holds
$\text{either}\quad\gamma(u,x_{0})=1\quad\text{or}\quad\gamma(u,x_{0})\geq 1+\delta_{d},$
for all $x_{0}\in\mathcal{F}(u)$, with $\delta_{d}>0$ being a dimensional constant. Once defined, for any $\gamma\geq 1$,
$\mathcal{F}_{\gamma}(u):=\{x\in\mathcal{F}(u)\colon\gamma(u,x_{0})=\gamma\},$
we split
$\mathcal{F}(u)=\mathrm{Reg}(u)\cup\mathrm{Sing}(u),$
where
$\mathrm{Reg}(u):=\mathcal{F}_{1}(u)\quad\text{and}\quad\mathrm{Sing}(u):=\bigcup_{\gamma\geq 1+\delta_{d}}\mathcal{F}_{\gamma}(u),$
with $\mathrm{Reg}(u)$ relatively open and $\mathrm{Sing}(u)$ relatively closed.
1.1.2. Blow-up analysis
@@
1.1.3. Points of frequency 1
Let $u\in\mathcal{S}(N,D)$ and let $x_{0}\in\mathcal{F}_{1}(u)$. Then, the limit $U\in\mathcal{S}_{1}(\mathbb{R}^{d};N)$ is unique (i.e. independent from the chosen subsequence) and of the form
$U_{1}=x_{d}^{-},\quad U_{2}=x_{d}^{+},\quad U_{i}\equiv 0~{}\text{for }i=3,\dots,N,$
up to rotations and dilations and there holds
$\frac{u(rx+x_{0})}{r}\to a\,U,~{}\text{uniformly in $B_{R}$ and in $H^{1}(B_{R};\mathbb{R}^{N})$, as }r\to 0,$
for some $a\neq 0$ and all $R>0$. Moreover, around $x_{0}$, the free interface $\mathcal{F}(u)$ is a smooth manifold separating two of the sets $\Omega_{i}^{u}$. Precisely, there is a neighborhood $B_{r}(x_{0})$ and two different indices $i,j\in\{1,\dots,N\}$ such that
$\Omega_{k}^{u}\cap B_{r}(x_{0})=\emptyset\quad\text{for every}\quad k\notin\{i,j\},$
while in $B_{r}(x_{0})$ the function $u_{i}-u_{j}$ is harmonic and
$|\nabla(u_{i}-u_{j})(x_{0})|\neq 0.$
In other words, by choosing $r$ small enough, the interface
$\{u_{i}-u_{j}=0\}\cap B_{r}(x_{0}),$
is smooth $(d-1)$-manifold and we have
$\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}^{u}\cap B_{r}(x_{0})=\partial\Omega_{j}^{u}\cap B_{r}(x_{0})=\{u_{i}-u_{j}=0\}\cap B_{r}(x_{0}).$
The same statements are true replacing $\mathcal{S}(D;N)$ with $\mathcal{M}(D;N)$ and $\mathcal{S}_{1}(\mathbb{R}^{d};N)$ with $\mathcal{M}_{1}(\mathbb{R}^{d};N)$ (which, actually, coincides with $\mathcal{S}_{1}(\mathbb{R}^{d};N)$).
1.1.4. Classification of blow-ups in dimension 2
In dimension $d=2$, the admissible frequencies and the blow-up limits are completely classified. Namely, it is known that if $x_{0}\in\mathcal{F}(u)$, then
$\gamma(u,x_{0})=\frac{m}{2}\quad\text{for some }m\in\mathbb{N},\ m\geq 2\,.$
Moreover, the sets $\mathcal{S}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ and $\mathcal{M}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ coincide and they contain a unique (up to rotations and dilations) segregated, $\nicefrac{{m}}{{2}}$-homogeneous function. More precisely, if we assume that all the $N$ components are non-trivial, then $N=m$ and any blow-up limit $u\in\mathcal{S}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)=\mathcal{M}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ is of the form
$u_{i}(r,\theta)=\begin{cases}r^{\frac{m}{2}}\left|\cos\left(\frac{m}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{m}(i-1)\leq\theta\leq\pi-\frac{2\pi}{m}i,\\ 0,&\text{elsewhere},\end{cases}$
up to rotations and dilations, for $i=1,\dots,N$.
1.1.5. Clean-up
In [Sun03], the author works in the framework of harmonic maps with $\mathbb{R}$-trees as target space (such as $\Sigma_{N}$). More precisely, in [Sun03, Theorem 1.2] it is proved what, in the language of optimal partition problems, is known as a clean-up result. Once translated in our framework, it says that for any $\gamma\geq 1$ there exists $\varepsilon,\rho>0$ depending on $d$ and $\gamma$ such that if $u\in\mathcal{M}(D;N)$ satisfies
$\mathcal{N}(u,r,x_{0})\leq\gamma\quad\text{and}\quad\left\lVert u_{i}\right\lVert_{L^{\infty}(B_{r}(x_{0}))}\leq\varepsilon\sqrt{H(u,r,x_{0})},$
for some $x_{0}\in D$, $r>0$ (such that $\overline{B_{r}(x_{0})}\subseteq D$) and $i\in I$, with $I\subseteq\{1,\dots,N\}$, then
$u_{i}\equiv 0\quad\text{in }B_{r\rho}(x_{0})\quad\text{for all }i\in I.$
In broad terms, if some components of a minimizer are small in a ball, then they disappear in a smaller ball.
1.1.6. The singular set
By a dimension reduction argument, for any $u\in\mathcal{M}(D;N)$ we have
$\dim_{\mathcal{H}}\mathrm{Sing}(u)\leq d-2.$
Furthermore, thanks to the techniques introduced in [NV17, NV18], in [Alp20, Dee22] it was shown that $\mathrm{Sing}(u)$ is $(d-2)$-rectifiable and that satisfies the following estimate
$|B_{r}(\mathrm{Sing}(u)\cap K)|\leq C(K,u)r^{2},$
for all compact $K\subseteq D$ and all $r>0$ sufficiently small.
1.2. New results on the structure of the singular set
Our main result (Theorem 1.1) provides a fine description of the singular set and of the whole nodal set around points in the lowest stratum. The first question we address is the detection of the optimal frequency gap $\delta_{d}$, defined in Section 1.1.1; until now, the frequency gap constant was explicitly known only in dimension $2$, i.e. $\delta_{2}=\nicefrac{{1}}{{2}}$ (see Section 1.1.4). Hence, our first main result extends this property to any dimension $d\geq 2$. We point out that this is true not only for local minimizers, but for all functions in the class $\mathcal{S}(D;N)$.
Theorem 1.2 (Optimal gap).
Let $u\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ be a non-trivial, $\gamma$-homogeneous critical point, with $\gamma\geq 1$. Then, either
$\gamma=1\quad\text{and}\quad N=2$
or
$\gamma\geq\frac{3}{2}\quad\text{and}\quad N\geq 3.$
Moreover, if $\gamma=\nicefrac{{3}}{{2}}$, then $N=3$ and, up to rotations and dilations, $u$ takes the form of a cylindrical $Y$-configuration, that is
$u(x_{1},\dots,x_{d})=Y(x_{d-1},x_{d}),$
where $Y=(Y_{1},Y_{2},Y_{3})\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{2};3)$ is defined as
for $i=1,2,3$.
The above theorem proves that for any $u\in\mathcal{S}(D,N)$ (and, in particular, for any minimizer $u\in\mathcal{M}(D,N)$) the singular set can be decomposed as follows
$\mathrm{Sing}(u)=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
Moreover, every blow-up limit of $u$ at a point $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ is of the form $Y$. This, of course, does not imply that the set
$\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}$
is smooth around $x_{0}$ and not even that the blow-up at $x_{0}$ is unique (as it might depend on the sequence of rescalings $u_{r_{n},x_{0}}$). We next focus on the regularity and the geometrical structure of set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$. Our approach in tackling this question is through an epiperimetric inequality for $\nicefrac{{3}}{{2}}$-Weiss energy, defined as
$W_{\frac{3}{2}}(u):=E(u,1,0)-\frac{3}{2}H(u,1,0)=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\frac{3}{2}\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S,$
for $u\in H^{1}(B_{1};\Sigma_{N})$. More precisely, we have the following.
Theorem 1.3 (Epiperimetric inequality).
and
$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
We prove Theorem 1.2 in Section 2. Then, in Section 4 and Section 5 we prove Theorem 1.3. Finally, using this epiperimetric inequality, in Section 6 we prove Theorem 1.1.
1.3. The Bishop-Friedland-Hayman’s conjecture for the min-max
In this section we use Theorem 1.2 to show that the optimal min-max $3$-partition of the sphere is given by the triple junction configuration in any dimension. For any $p\in(0,+\infty]$ we consider the following family of optimal partition problems studied by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]
($\text{OP}_{p}$) $\min\left\{\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
where $\lambda_{1}(\omega)$ denotes the first Dirichlet eigenvalue of the spherical Laplacian in $\omega\subseteq\partial B_{1}$, and where in the limit case $p=+\infty$, the value
$\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}$
is intended as
$\max_{i\in\{1,\dots,N\}}\lambda_{1}(\omega_{i}).$
In the particular case of the $3$-partition, i.e. $N=3$, we have the following conjecture.
Problem 1.4 (Bishop-Friedland-Hayman-Helffer-Hoffmann-Ostenhof-Terracini).
For $N=3$, prove that the $Y$-configuration is the unique minimizer of ($\text{OP}_{p}$), for any $p\in(0,+\infty]$.
This conjecture was originally posed in [Bis92] in the case $p=1$, but can be generalized for any $p\in(0,+\infty]$ (see [FH76] and [HHT10]). Of particular interest is the case $p=1/2$, which is related with the best exponent in the multi-phase Alt-Caffarelli-Friedman monotonicity formula; in general, there is a vast literature on these optimal partition problems also in relation to models in biology and dynamics of populations, we refer for instance to [CTV03, CTV05b, BBH98, CL07, HHT09, HHT10]. 1.4 was proved for $p=+\infty$ in dimension $d=3$ by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]; in Section 3 we prove the following theorem, which solves 1.4 (for $p=+\infty$) in any dimension $d\geq 3$.
Theorem 1.5.
The $Y$-configuration is the unique minimizer of
$\min\left\{\max_{i\in\{1,2,3\}}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\}.$
1.4. Two consequences of Theorem 1.1
In this section we discuss two corollaries of Theorem 1.1 for harmonic maps with values in metric graphs and for spectral optimal partitions.
First of all we notice that an immediate corollary of Theorem 1.1 (and the clean-up theorem from [Sun03]), is that the same result holds for harmonic maps with values in a locally finite tree $\Sigma$. Indeed, by [GS92, Sun03, Dee22], if $u:B_{1}\to\Sigma$ is an energy-minimizing map defined in $B_{1}\subset\mathbb{R}^{d}$ with values in a locally finite metric tree, and if $p$ is a vertex of $\Sigma$, then the set $\mathcal{F}:=u^{-1}(\{p\})$ can be decomposed into a regular and a singular parts, $Reg$ and $Sing$, where $Reg$ is a a smooth $(d-1)$-manifold that consists of the points of order (frequency) ${\rm Ord}^{u}$, while $Sing$ has Hausdorff dimension $(d-2)$ and consists of of points of order (frequency) ${\rm Ord}^{u}>1$; furthermore, for every $x\in\mathcal{F}$, there is a neighborhood $\Omega\subset B_{1}$ such that $u:\Omega\to\Sigma$ has values in a subtree $\Sigma_{N}$ of the form (1.1). Thus, we have the following
Theorem 1.6.
Let $\Sigma$ be a metric tree which is locally finite and let $u:B_{1}\to\Sigma$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$. Then, for every vertex $p$ of $\Sigma$ the singular set of $\mathcal{F}:=u^{-1}(\{p\})$ can be decomposed as the disjoint union of
$\mathrm{Sing}=\Big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)>\nicefrac{{3}}{{2}}\Big{\}}.$
The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}:=\big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$, the set $\mathcal{F}$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}$.
Another consequence concerns the regularity of the following optimal partition problem studied in [CL07], [CTV03] and [OV24]. Given a bounded open set $D\subset\mathbb{R}^{d}$, $d\geq 2$, and a natural number $N\geq 2$ we consider the optimal partition problem
$\mathcal{F}:=D\cap\Big{(}\cup_{j}\partial\Omega_{j}\Big{)}$
can be decomposed as a regular part and a singular part as follows (see [CL07, Alp20]): the regular part $\mathrm{Reg}$ consists of all points $x\in\mathcal{F}$ of frequency $\gamma(u,x)=1$ and is a smooth manifold of dimension $d-1$; the singular part $\mathrm{Sing}$ is a closed $(d-2)$-rectifiable set composed of all the points of frequency $\gamma(u,x)>1$. We have the following theorem
Theorem 1.7.
Let $(\Omega_{1},\dots,\Omega_{N})$ be a minimizer of (1.2) and $u=(u_{1},\dots,u_{N})$ be the corresponding vector of the normalized first eigenfunctions. Then, the singular set of $\mathrm{Sing}$ of the free interface $\mathcal{F}$ can be decomposed as the disjoint union of
$\mathrm{Sing}=\Big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}:=\big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$, the set $\mathcal{F}$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}$.
2. Optimal gap
@@ 2.1. A topological lemma
Let $\mathcal{M}$ be a smooth $m$-dimensional manifold. We recall that a subset $\mathcal{F}\subset\mathcal{M}$ is a $C^{1}$ submanifold of dimension $m-1$ if for every $x_{0}\in\mathcal{F}$ there is an open neighborhood $\omega\subset\mathcal{M}$ of $x_{0}$ and a diffeomorphism $\Phi:\omega\to B_{1}\times(-1,1)$ such that $\Phi(\mathcal{F})$ is the graph $\Gamma_{\eta}$ of a $C^{1}$ function $\eta:B_{1}\to(-1,1)$; here $B_{1}$ is the unit ball in $\mathbb{R}^{m-1}$.
@@
Let $\mathcal{M}$ be a smooth $m$-dimensional manifold and let $\mathcal{F}\subset\mathcal{M}$ be a $C^{1}$ submanifold of dimension $m-1$. Given $x\in\mathcal{M}\setminus\mathcal{F}$ and $y\in\mathcal{M}\setminus\mathcal{F}$, we denote by $C(x\to y)$ the family of piecewise $C^{1}$ curves transversal to $\mathcal{F}$. Precisely, we say that $\ell\in\mathcal{C}(x\to y)$ if:
@@
    $Q_{i,j}$ is contained in an open set $\omega_{2}\subset\mathcal{M}$ for which there is a diffeomorphism $\Phi:\omega_{2}\to B_{1}\times(-1,1)$ such that $\Phi(\mathcal{F})$ is the graph $\Gamma_{\eta}$ of a $C^{1}$ function $\eta:B_{1}\to(-1,1)$, $B_{1}$ being the unit ball in $\mathbb{R}^{m-1}$.
@@
2.2. Proof of Theorem 1.2
In this section we prove Theorem 1.2. The proof is based on a dimension reduction argument and on the following

### v2

@@ 2.1. Lipschitz continuity
are open and well-defined and that $u_{i}$ are harmonic on $\Omega_{i}^{u}$.
2.2. Almgren and Weiss monotonicity formulas
For any non-trivial $u\in\mathcal{S}(D;N)$ one can define the Almgren frequency function for any $x_{0}\in D$ and any $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$ as
$\mathcal{N}(u,x_{0},r):=\frac{E(u,x_{0},r)}{H(u,x_{0},r)},$
@@ 2.2. Almgren and Weiss monotonicity formulas
$E(u,x_{0},r):=\frac{1}{r^{d-2}}E(u,B_{r}(x_{0}))=\frac{1}{r^{d-2}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}|\nabla u_{i}|^{2}\,\mathrm{d}x$
@@
$H(u,x_{0},r):=\frac{1}{r^{d-1}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}u_{i}^{2}\,\mathrm{d}S$
is the scaled height function. It is known that for $u\in\mathcal{S}(D;N)$ the function $r\mapsto\mathcal{N}(u,x_{0},r)$ is monotone non-decreasing and so the frequency of $u$ at $x_{0}$
$\gamma(u,x_{0}):=\lim_{r\to 0^{+}}\mathcal{N}(u,x_{0},r)$
is well-defined at any point $x_{0}\in\mathcal{F}(u)$. As a consequence, one can derive unique continuation properties implying that $\mathcal{F}(u)$ has empty interior and that
$\mathcal{F}(u)=\bigcup_{i=1}^{N}\partial\{x\in D\colon u_{i}(x)>0\}.$
Hence, we can also split the points of the free boundary in terms of their frequency; namely, for any $\gamma>0$ we let
$\mathcal{F}_{\gamma}(u):=\{x\in\mathcal{F}(u)\colon\gamma(u,x)=\gamma\}.$
Finally, we point out that also the Weiss energy
$W_{\gamma}(u,x_{0},r):=\frac{H(u,x_{0},r)}{r^{2\gamma}}\Big{(}\mathcal{N}(u,x_{0},r)-\gamma\Big{)}$
is monotone non-decreasing with respect to $r>0$, for any $\gamma>0$.
2.3. Blow-up analysis
@@
3. Singular blow-ups of minimal frequency
The aim of this section is to classify homogeneous singular solutions with minimal frequency $\nicefrac{{3}}{{2}}$, i.e. functions in $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$.
@@ 3.1. A topological lemma
This section is dedicated to a topological lemma, which we use both in the classification of the $\nicefrac{{3}}{{2}}$-homogeneous blow-ups (Proposition 3.5) and in the proof of the no-holes lemma (Lemma 6.6).
@@
Let $\mathcal{M}$ be a smooth $m$-dimensional manifold and let $\mathcal{F}\subset\mathcal{M}$ be a $C^{1}$ submanifold of dimension $m-1$ (that is, for every $x_{0}\in\mathcal{F}$ there is an open neighborhood $\omega\subset\mathcal{M}$ of $x_{0}$ and a $C^{1}$-diffeomorphism $\Phi:\omega\to B_{1}\subset\mathbb{R}^{d}$ such that $\Phi(\mathcal{F})=B_{1}\cap\{x_{m}=0\}$). For any $x,y\in\mathcal{M}\setminus\mathcal{F}$, we denote by $C(x\to y)$ the family of piecewise $C^{1}$ curves transversal to $\mathcal{F}$. Precisely, we say that $\ell\in\mathcal{C}(x\to y)$ if:
@@
    $Q_{i,j}$ is contained in an open set $\omega_{2}\subset\mathcal{M}$ for which there is a $C^{1}$-diffeomorphism $\Phi:\omega_{2}\to B_{1}\subset\mathbb{R}^{m}$ such that $\Phi(\mathcal{F})=B_{1}\cap\{x_{m}=0\}$.
@@
3.2. Classification of the blow-ups of frequency $3/2$
In this section we prove that any $\nicefrac{{3}}{{2}}$-homogeneous function $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$ is of the form $Y$ (see Proposition 3.5). We stress that even if this result is not explicitly stated in [ST15], the key elements (the dimension reduction argument and Lemma 3.4) are already contained in the proof of [ST15, Lemma 4.2]; here below we give the details since we need Proposition 3.5 in Theorem 1.1.